OpenStudy (lgbasallote):

A particle moves along a circular path having a radius of 2 m. At an instant when the speed of the particle is equal to 3 m/s and changing at the rate of 5 m/s^2, what is the magnitude of the total acceleration of the particle?

OpenStudy (lgbasallote):

first off...what does total acceleration mean?

OpenStudy (ghazi):

i guess motion is on a vertical plane ....that might be including acceleration due to gravity i am not sure though

OpenStudy (lgbasallote):

im thinking radial acceleration and tangential acceleration?

OpenStudy (lgbasallote):

this is horizontal circle btw...

OpenStudy (lgbasallote):

i think

OpenStudy (ghazi):

then it is fine

mathslover (mathslover):

Should I interrupt? or Did I :P ?

OpenStudy (ghazi):

@mathslover go on mate

mathslover (mathslover):

thanks... so my first suggestion is : it is related to centripetal acceleration

OpenStudy (lgbasallote):

yes...

mathslover (mathslover):

So what will be centripetal acc.

OpenStudy (lgbasallote):

which is \[\frac{v^2}r\] yes?

OpenStudy (ghazi):

\[a= \frac{ v^2 }{ r }\]

OpenStudy (ghazi):

i guess it is \[toatal ..a=\sqrt{(4.5)^2+(5)^2}\]

OpenStudy (lgbasallote):

ahh yes that's right

OpenStudy (ghazi):

6.72 is it?

OpenStudy (lgbasallote):

out of curiousity...what is the formula for the other kind of acceleration?

OpenStudy (lgbasallote):

i know there are two kinds of acceleration associated with circle...

OpenStudy (ghazi):

you just need to find the resultant acceleration one is tangential and one is centripetal

OpenStudy (ghazi):

tangential stands for linear motion at that point for which we have three laws in kinematics and for centripetal we have v^2/r

OpenStudy (lgbasallote):

oh okay thanks

OpenStudy (lgbasallote):

is there a formula for tangential acceleration?

OpenStudy (ghazi):

three laws of kinematics

OpenStudy (ghazi):

resultant acceleration =\[\sqrt{(\frac{ v^2 }{ r })^2+(tangetial)^2}\]

OpenStudy (lgbasallote):

what are the three laws of kinematics?

OpenStudy (ghazi):

v=u+a*t v^2=u^2+at s=u*t+0.5*a*t^2

OpenStudy (ghazi):

*2*a*s

OpenStudy (ghazi):

in the second equation

OpenStudy (lgbasallote):

ohhh that

OpenStudy (lgbasallote):

that's tangential acceleration? o.O

OpenStudy (ghazi):

tangential acceleration is the acceleration of a body at any instant